OpenStudy (anonymous):

HELP PLEASE!!!!! Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20 cospi over 15 + 30, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

Part 1: How far above the ground is Xavier before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum? You must show all work

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@nikato

OpenStudy (anonymous):

@Cosmichaotic

OpenStudy (anonymous):

@IMStuck

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@MaimiGirl

OpenStudy (triciaal):

part 1 is asking what is h when t = 0

OpenStudy (anonymous):

so 30 right?

OpenStudy (triciaal):

are you sure you have the eqn written correctly. H (t) but no t in the expression 20 cos pi/15 = 0

OpenStudy (anonymous):

Sorry. H(t)=20cos(pi/15)t+30

OpenStudy (triciaal):

ok now you have the funnction when t = 0 h = 30 yes

OpenStudy (anonymous):

what about part 2 and 3

OpenStudy (triciaal):

find what is the value of t when h = 0.

OpenStudy (anonymous):

how do you do that

OpenStudy (triciaal):

set the function =0

OpenStudy (triciaal):

H(t) =( 20 cospi over 15) t+ 30 = 0

OpenStudy (anonymous):

So t=1.5

OpenStudy (anonymous):

sorry -1.5

OpenStudy (triciaal):

let me try to get some more input . not very fast with this anymore.

OpenStudy (triciaal):

@Kainui can you help with this?

OpenStudy (anonymous):

can you help @Kainui

OpenStudy (anonymous):

ANYONE!!!!!!

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

let @Kainui help. He is better than me, hihihi

OpenStudy (anonymous):

when t =0 H(t) = 50 because cos 0 =1

OpenStudy (anonymous):

the equation of the height is H(t) =\(20cos(\dfrac{\pi t}{15})+30\) right?

OpenStudy (anonymous):

confirm or correct this equation, please

OpenStudy (anonymous):

Anyway, I think the net is down so that you can't reply for b) take derivative and let it =0, solve for t, (the reason is finding the max and min of the graph) then times that t by2 to get the answer

OpenStudy (anonymous):



OpenStudy (anonymous):

for c) the min is one of the roots we get from b) replace the value of min into the H(t) to get the height at min, and then at max, heigh max-heigh min= answer